### v2

@@ 3.2 DYNAMIC SENSOR PLACEMENT
Note that (24) is based on a greedy selection introduced in Section 22.2. However, in practice, this optimization may not be efficient for large graphs: The control law with (24) implies that we need a large computational burden if we determine the positions of all sensors. This results in a delay for sensor relocations which should be alleviated. In this paper, instead, we consider selecting the sensor positions independently, i.e., (24) is solved for a sensor independently of the other ones. In the following, we rewrite (24) as a distributed optimization.
@@
Input: Sensing matrix $\mathbf{G}$, initial positions of sensors $\mathcal{Q}_{1}$, time duration $T$
@@
       Update the data matrix $\tilde{\mathbf{X}}_{t}$
@@
where $d(v,q_{i,t})$ is the shortest distance between $v$ and $q_{i,t}$, and $\mathcal{Q}_{t}=\{q_{j,t}\}_{j=1,\ldots,K}$ denotes the set of all sensor positions at $t$. Note that, in each $\mathcal{W}_{i,t}$, only one sensor is contained: We illustrate its example in Fig. 2.
@@
$q^{*}_{i,t+1}=\mathop{\rm arg~{}max}\limits_{y\in\mathcal{W}_{i,t}\cap\mathcal{N}_{i,t}^{P}}\|\bm{\nu}_{y}\|^{2}-\|\mathbf{P}_{\mathcal{R}(\mathbf{N}_{:\mathcal{M}_{i,t}})}\bm{\nu}_{y}\|^{2}.$ (28)
Algorithm 2 summarizes the proposed dynamic sensor placement method [^3] We utilize the MATLAB notation where we denote the parallel processing in Algorithm 2 as par for..
Note that our main contribution is to formulate dynamic sensor placement as graph signal sampling. We may be able to use other methods as an internal algorithm for dictionary learning like [39] and [40].
@@ 4 EXPERIMENTS
We suppose that the graph $\mathcal{G}$ is given a priori. The nodes represent the candidates of sensor positions, and the edges indicate the paths where sensors can move. Sampling is performed at every time instance. After sampling, the reconstructed signal at $t$, i.e., $\tilde{\mathbf{x}}_{t}$, is stored as the latest signal collection $\tilde{\mathbf{X}}_{t-1}$ which is used for the dictionary/sensor positions update.
Figure 4: Reconstructed graph signals for PC signals at time $t=4,9,13$. Circled nodes represent selected sensor positions.
@@ 4.1.1 SETUP
For both graph signal models, we generate the data matrix $\mathbf{X}_{t}$ with $T=40$. The number of time slots is set to $D=20$.
While the initial expansion coefficients and dictionary can be arbitrary as mentioned in Section 33.1. In the experiment, however, we used an SVD-based dictionary for a fair comparison to alternative methods. Therefore, $\mathbf{A}_{0}=\bm{\mathcal{U}}$ at $t=0$, where $\bm{\mathcal{U}}$ is obtained by the SVD of the known measurement data, i.e., $\mathbf{X}_{0}=\bm{\mathcal{U}}\mathbf{\Sigma}\bm{\mathcal{V}}^{\top}$. The initial value of $\mathbf{D}_{t}$ is also set to $\mathbf{D}_{0}=\mathbf{11}^{\top}$. We set the temporal sampling period $T_{s}=\frac{\pi}{30}$ to avoid aliasing. The parameters of the dictionary learning are experimentally set to $(\gamma,\eta,\mu)=(10^{-4},3,1)$, and $P=1$ is used for the sensor movable area where we assume a sensor can only be moved to its one-hop neighbor based on reasonable sensor mobility.
@@
Recovery accuracy is measured in noisy cases where white Gaussian noise $w[i]\sim\mathcal{N}(0,0.1)$ is added to the signals. For all methods, the initial position of sensors is determined by a state-of-the-art sampling set selection method on graphs [8]. We calculate the averaged MSEs for 50 independent runs.
@@ 4.1.2 RESULTS
From Fig. 3, the proposed method shows consistently (and significantly) smaller MSEs than the other methods in all time instances, and the MSE of the proposed method gradually decreases over time. This indicates that the proposed method can adapt to the change in measurement conditions, while the static methods fail to do so.
In this experiment, $P=1$ and $\infty$ do not results in a large difference. However, as shown in Fig. 4 at $t=13$, we observe that $P=\infty$ produces a larger error than $P=1$ because several sensor locations are very close. Its possible reason is that the sensor positions were biased on the right side.
@@ 4.2.1 SETUP
We use the global sea surface temperature data [41] during 2016–2021. This dataset is composed of snapshots recorded every month from 2016 to 2021, i.e., its duration is $T=60$. Temperature is sampled at intersections of 1-degree latitude-longitude grids. For simplicity, we clip snapshots of the west coast of the U.S. and randomly sample 100 nodes for the experiments. Then, we construct a $5$-nearest neighbor graph. The numbers of time slots and sensors are set to $D=5$ and $K=10$, respectively. The parameters of the dictionary learning are set to $(\gamma,\eta,\mu)=(10^{-3},1,1)$, and $P=1$ is used for the sensor movable area. We design the filter matrix $\mathbf{G}$ the same as (31) in Section 44.1. We evaluate the averaged MSEs and compare them with the above-mentioned two benchmark methods Static1 and 2.

### v4

@@ 3.2 Dynamic Sensor Placement
Note that (25) is based on a greedy selection introduced in Section 2-2.2. However, in practice, this optimization may not be efficient for large graphs: The control law with (25) implies that we need a large computational burden if all sensors’ positions are sequentially determined at every time instance. This results in a delay for sensor relocations which should be alleviated. In this paper, instead, we consider selecting the sensor positions independently, i.e., (25) is solved for a sensor independently of the other ones. In the following, we rewrite (25) as a distributed optimization.
@@
Input: Sensing matrix $\mathbf{G}$, time duration $T$, initial data slot $\mathbf{X}_{0}$, initial dictionary $\mathbf{A}_{0}$, initial expansion coefficients matrix $\mathbf{D}_{0}$, initial positions of sensors $\mathcal{Q}_{1}$
@@
       Sample the current signal $\mathbf{x}_{t}$ at the positions $\mathcal{Q}_{t}$ in (5)
       Reconstruct the current signal $\tilde{\mathbf{x}}_{t}$ in (6)
       Update the data matrix $\tilde{\mathbf{X}}_{t}$ by inserting $\tilde{\mathbf{x}}_{t}$
@@
where $d(v,q_{i,t})$ is the shortest distance between $v$ and $q_{i,t}$, and $\mathcal{Q}_{t}=\{q_{j,t}\}_{j=1,\ldots,K}$ denotes the set of all sensor positions at $t$. Note that, in each $\mathcal{W}_{i,t}$, only one sensor is contained: We illustrate its example in Fig. 2. The Voronoi regions are determined by $k$-means clustering, which is robust to perturbations to some extent depending on the topology [38]. Detailed analysis of robustness is left for future work.
@@
$q^{*}_{i,t+1}=\mathop{\rm arg~{}max}\limits_{y\in\mathcal{W}_{i,t}\cap\mathcal{N}_{i,t}^{P}}\Psi(y),$ (29)
where $\Psi(y):=\|\bm{\nu}_{y}\|^{2}-\|\mathbf{P}_{\mathcal{R}(\mathbf{N}_{:\mathcal{M}_{i,t}})}\bm{\nu}_{y}\|^{2}$. Algorithm 2 summarizes the proposed dynamic sensor placement method [^2] We utilize the MATLAB notation where we denote the parallel processing in Algorithm 2 as par for..
As mentioned in Section 2-2.2, we utilize the D-optimality design for sensor selection, which is a monotone submodular function. It is widely known that greedy sensor selection with monotone submodular functions yields a nearly optimal solution [39].
Note that we consider a distributed sensor placement problem, which is an approximation of the greedy (centralized) one in some sense. We address the gap between the distributed and centralized solutions. Here, we denote the set of sensor positions obtained in a distributed manner by $\{q_{i}^{*}\}_{i=1,\ldots,K}$ and those in a centralized manner by $\{r_{i}^{*}\}_{i=1,\ldots,K}$. From [40, Theorem 4.1], we have $\sum_{i=1}^{K}\Psi(r_{i}^{*})\geq\sum_{i=1}^{K}\Psi(q_{i}^{*})\geq\frac{1}{K}\sum_{i=1}^{K}\Psi(r_{i}^{*})$.
Note that our main contribution is to formulate dynamic sensor placement as graph signal sampling. We may be able to use other methods, such as those described in [41] and [42], as internal algorithms for dictionary learning.
3.3 Computational Complexity
We now address the computational complexity of the proposed sensor placement method. Determining the sensor placement in (11) requires $O(K^{3}+NK^{2})$ computational complexity, where $O(K^{3})$ complexity costs for computing the inverse in (11), and $O(NK^{2})$ complexity requires for the greedy search of the optimal nodes. While (11) determines the positions from the whole network [18], our algorithm performs it within the Voronoi regions in the distributed manner. As a result, the computational complexity of our algorithm results in $O(K(N+|\mathcal{E}|)+K^{3}+\max_{i}\{|\mathcal{W}_{i,t}|\})$ at most, where $O(K(N+|\mathcal{E}|))$ complexity requires for the division of Voroinoi regions and $O(\max_{i}\{|\mathcal{W}_{i,t}|\})$ complexity is needed for the node selection in the Voronoi reigions in the worst case. By comparing $O(NK^{2})$ and $O(K(N+|\mathcal{E}|)+\max_{i}\{|\mathcal{W}_{i,t}|\})$ complexities, the latter is usually lower as long as the graph is not very dense.
@@ 4 EXPERIMENTS
We suppose that the graph $\mathcal{G}$ is given a priori. The nodes represent the candidates of sensor positions, and the edges indicate the paths where sensors can move. Sampling is performed at every time instance. After sampling, the reconstructed signal at $t$, i.e., $\tilde{\mathbf{x}}_{t}$, is stored as the latest signal collection $\tilde{\mathbf{X}}_{t}$, which is used for the dictionary/sensor positions update. Simultaneously, the oldest signal, $\tilde{\mathbf{x}}_{t-D}$, is discarded from $\tilde{\mathbf{X}}_{t}$.
@@ 4.1.1 Setup
Figure 4: Reconstructed graph signals for PC signals at time $t=6,10,14$. Circled nodes represent selected sensor positions.
@@
For both graph signal models, we generate the graph signals $\mathbf{x}_{t}$ for durations of $T_{\text{train}}=20$ for training and $T_{\text{test}}=20$ for test, respectively. The number of time slots is set to $D=20$.
The initial data slot $\mathbf{X}_{0}$ used for the dictionary learning in (16) is set to $\mathbf{X}_{0}=\mathbf{X}_{\text{train}}+\mathbf{M}$, where $\mathbf{X}_{\text{train}}$ is the collection of original signals for training, and $\mathbf{M}_{i,j}\sim\mathcal{N}(0,0.1)$ is white Gaussian noise. The initial expansion coefficients and dictionary can be arbitrary as mentioned in Section 3-3.1. In the experiment, however, we used an SVD-based dictionary for a fair comparison to alternative methods. Therefore, $\mathbf{A}_{0}=\bm{\mathcal{U}}$ at $t=0$, where $\bm{\mathcal{U}}$ is obtained by the SVD of the known measurement data, i.e., $\mathbf{X}_{0}=\bm{\mathcal{U}}\mathbf{\Sigma}\bm{\mathcal{V}}^{\top}$. The initial value of $\mathbf{D}_{t}$ is also set to $\mathbf{D}_{0}=\mathbf{11}^{\top}$. We set the temporal sampling period $T_{s}=\frac{\pi}{30}$ to avoid aliasing. The parameters of the dictionary learning are experimentally set to $(\gamma,\eta,\mu)=(10^{-4},3,1)$, and $P=1$ is used for the sensor movable area where we assume a sensor can only be moved to its one-hop neighbor based on reasonable sensor mobility.
@@
We apply the Chebyshev polynomial approximation [43] to $\mathbf{G}$. This works as a $J$-hop localized filter. We set the order of approximation to $J=20$.
Recovery accuracy is measured in noisy scenarios where white Gaussian noise $\mathbf{m}_{i}\sim\mathcal{N}(0,0.1)$ is added to the signals (see (5)). For all methods, the initial position of sensors is determined by a state-of-the-art sampling set selection method on graphs [8]. We calculate the averaged mean squared errors (MSEs) for 50 independent runs.
@@ 4.1.2 Results
Fig. 3 indicates that the proposed method shows consistently (and significantly) smaller MSEs than the other methods in all time instances, and the MSE of the proposed method gradually decreases over time. This validates that the proposed method can adapt to the change in measurement conditions, while the static methods fail to do so.
In this experiment, $P=1$ and $\infty$ do not result in a large difference. However, as shown in Fig. 4 at $t=14$, we observe that $P=\infty$ produces a larger error than $P=1$ because several sensor locations are very close to each other. A possible reason for this is that the sensor positions were biased toward the right side.
@@ 4.2.1 Setup
We use the global sea surface temperature data [44] during 2016–2021. This dataset is composed of snapshots recorded every month from 2016 to 2021. The first five months are used for training ($T_{\text{train}}=5$), while the remaining months are used for testing ($T_{\text{test}}=55$). The initial data slot $\mathbf{X}_{0}$ of the dictionary learning is set to the same scenario as in the synthetic experiments. Temperature is sampled at intersections of 1-degree latitude-longitude grids. For simplicity, we clip snapshots of the west coast of the U.S. and randomly sample 100 nodes for the experiments. Then, we construct a $5$-nearest neighbor graph. The numbers of time slots and sensors are set to $D=5$ and $K=10$, respectively. The parameters of the dictionary learning are set to $(\gamma,\eta,\mu)=(10^{-3},1,1)$, and $P=1$ is used for the sensor movable area. We design the filter matrix $\mathbf{G}$ the same as (32) in Section 4-4.1. We evaluate the averaged MSEs and compare them with the above-mentioned two benchmark methods Static1 and 2.
Figure 7: Comparison of the averaged MSE of the proposed method with different numbers of sensors $K$.
@@ 4.2.2 Results
We also compare the averaged MSE of the proposed method with varying numbers of sensors (different values of $K$). Fig. 7 shows the experimental results with different $K$. As expected, the proposed method tends to have a smaller average MSE for a larger $K$.